# Unified Source-Free Domain Adaptation

### v1

@@ 5 Conclusion
This paper considers a more challenging Unified SFDA problem that thoroughly addresses all specific scenarios in a unified manner. Further, we propose a LCFD approach from a causality view, conceptually different from prior statistical dependence-based scheme. Specifically, we propose a self-supervised information bottleneck to capture the external ones and further transform them into pseudo-labels to regulate the successive IC process for discovering internal ones. According to our theoretical result, we instantiate the proposed information bottleneck based on the network implementation for variational mutual information. Extensive experiments on four mainstream SFDA settings and SF-OODG validate the superiority of our LCFD.
@@ Acknowledgments
This work is partially supported by the German Research Foundation and National Natural Science Foundation of China (NSFC) in project Crossmodal Learning under contract Sonderforschungsbereich Transregio 169, the Hamburg Landesforschungsförderungsprojekt Cross, NSFC (61773083); Horizon2020 RISE project STEP2DYNA (691154); NSFC (62206168, 62276048).

### v4

@@ 5 Conclusion
We introduce a more challenging Unified SFDA problem that comprehensively addresses all specific scenarios within a unified framework. To tackle this, we propose CausalDA, a novel approach grounded in causality, fundamentally differing from previous methods based on statistical dependence. Specifically, we design a self-supervised information bottleneck to capture external causal elements, convert them into pseudo-labels, and use them to guide the discovery of internal causal factors. Building on our theoretical insights, we implement this information bottleneck using a network-based variational mutual information approach. Extensive experiments across various SFDA settings and SF-OODG validate the superiority of CausalDA over existing methods.
@@ Acknowledgments
This work is partially supported by the National Natural Science Foundation of China (NSFC) (62476169, 62206168, 62276048); the German Research Foundation and NSFC in project Crossmodal Learning under contract Sonderforschungsbereich Transregio 169, the Hamburg Landesforschungsförderungsprojekt Cross, NSFC (61773083); Horizon2020 RISE project STEP2DYNA (691154).
Appendix A A Proof of Lemma 1
Restatement of Lemma 1 Given random variables $Z_{1}$, $X_{1}$, $Y_{1}$ where $X_{1}$, $Y_{1}$ satisfy a mapping $f_{1}\!:\!X_{1}\!\mapsto\!Y_{1}$. When $f_{1}$ is compressed, i.e., the output’s dimension is smaller than the input’s,
$\small\begin{split}I\left({{Z_{1}},X_{1}}\right)\geq I\left({{Z_{1}},Y_{1}}\right)\end{split}$ (19)
Proof. Since mapping $f_{1}$ is compressed, the entropy of $X_{1}$, $Y_{1}$ have $H(X_{1})\geq H(Y_{1})$. As for $I\left({{Z_{1}},X_{1}}\right)$, it has
$\small\begin{split}I\left({{Z_{1}},X_{1}}\right)&\geq I\left({{Z_{1}},X_{1}}\right)-H\left(Z_{1}|Y_{1}\right)-I\left(X_{1},Y_{1}\right)\\ &\geq\left[H\left(Z_{1}\right)+H\left(X_{1}\right)-H\left(Z_{1},X_{1}\right)\right]-H\left(Z_{1}|Y_{1}\right)\\ &~~~~~-\left[H\left(X_{1}\right)+H\left(Y_{1}\right)-H\left(X_{1},Y_{1}\right)\right]\\ &=H\left(Z_{1}\right)-H\left(Z_{1}|Y_{1}\right)-H\left(Y_{1}\right)+H\left(X_{1},Y_{1}\right)\\ &\geq I\left({{Z_{1}},Y_{1}}\right)-H\left(Y_{1}\right)+\max\left(H\left(X_{1}\right),H\left(Y_{1}\right)\right)\\ &\geq I\left({{Z_{1}},Y_{1}}\right)-H\left(Y_{1}\right)+H\left(X_{1}\right)\\ &\geq I\left({{Z_{1}},Y_{1}}\right)-H\left(Y_{1}\right)+H\left(Y_{1}\right)\\ &=I\left({{Z_{1}},Y_{1}}\right)\end{split}$ (20)
Appendix B A Proof of Theorem 1
Restatement of Theorem 1 Suppose there are five random variables, $Z$, $V$, $Z{{}^{\prime}}$, $Y$ and $Y{{}^{\prime}}$. Among them, $Z$ represents the target domain knowledge; $V$, $Z{{}^{\prime}}$ and $Y$ express the input instances, predictions and intermediate features of ViL model, respectively; $Y{{}^{\prime}}$ depicts a pseudo-label that has a learnable functional relationship $f_{w}$ ($w$ is parameters) with $Z{{}^{\prime}}$. When the function $f_{w}$ is reversible and uncompressed, we have this upper bound as follows.
$\small\begin{split}&\min_{\boldsymbol{v}_{d}}I\left({Z,{Z{{}^{\prime}}}}\right)-I\left({{Z{{}^{\prime}}},Y}\right)\\ &\leq\min_{\boldsymbol{v}_{d},w}\underbrace{I\left({Z,{Z{{}^{\prime}}}}\right)}_{{T_{1}}}-\underbrace{I\left({{Y{{}^{\prime}}},Y}\right),{Y{{}^{\prime}}}=f_{w}\left({{Z{{}^{\prime}}}}\right)}_{{T_{2}}}.\end{split}$ (21)
Proof. Since the function $f_{w}$ is reversible and uncompressed, the $Z{{}^{\prime}}$ and $Y{{}^{\prime}}$ satisfy bidirectional one-to-one mapping. In other words, the from $Y{{}^{\prime}}$ to $Z{{}^{\prime}}$, there is not information loss, formally $H\left(Y{{}^{\prime}}\right)=H\left(Z{{}^{\prime}}\right)$, termed condition A. Let $L=I\left({Z,Z{{}^{\prime}}}\right)-I\left({Z{{}^{\prime}},Y}\right)$, we have
$\small\begin{split}L&\leq I\left({Z,Z{{}^{\prime}}}\right)-I\left({Z{{}^{\prime}},Y}\right)+I\left({Z{{}^{\prime}},Y{{}^{\prime}}}\right)\\ &\leq I\left({Z,Z{{}^{\prime}}}\right)-\left[H\left({Z^{{}^{\prime}}|Y{{}^{\prime}}}\right)-H\left({Z{{}^{\prime}}|Y}\right)\right].\end{split}$ (22)
The following discusses separately the two components in the second item of Eq. (22) under condition A.
(1) Proof $H\left({Z{{}^{\prime}}|Y{{}^{\prime}}}\right)=H(Y^{{}^{\prime}})$.
Let $L_{1}=H\left({Z{{}^{\prime}}|Y{{}^{\prime}}}\right)$, we have the relation as
$\small\begin{split}L_{1}&=H\left({Z{{}^{\prime}},Y{{}^{\prime}}}\right)-H\left(Y{{}^{\prime}}\right)\\ &\leq H\left(Z{{}^{\prime}}\right)+H\left(Y{{}^{\prime}}\right)-H\left(Y{{}^{\prime}}\right)=H\left(Z{{}^{\prime}}\right)\end{split}$ (23)
Also, we know $L_{1}\geq H\left(Z{{}^{\prime}}\right)$. Combing Eq. (23), we obtain $L_{1}=H\left(Z{{}^{\prime}}\right)=H\left(Y{{}^{\prime}}\right)$ according to Sandwich Theorem and condition A.
(2) Proof $H\left({Z{{}^{\prime}}|Y}\right)=H\left({Y{{}^{\prime}}|Y}\right)$.
Let $L_{2}=H\left({Z{{}^{\prime}}|Y}\right)$, $L_{3}=H\left({Y{{}^{\prime}}|Y}\right)$, here, we prove $L_{2}=L_{3}$ with condition A. For $L_{2}$, it can be expressed as
$\small L_{2}=H\left(Z{{}^{\prime}},Y\right)-H\left(Y\right),$ (24)
then we have
$\small\begin{split}\max\left(H\left(Z{{}^{\prime}}\right),H\left(Y\right)\right)-H\left(Y\right)&\leq L_{2}\leq H\left(Z{{}^{\prime}}\right)\\ H\left(Z{{}^{\prime}}\right)-H\left(Y\right)&\leq L_{2}\leq H\left(Z{{}^{\prime}}\right).\end{split}$ (25)
Similarly, $L_{3}$ can be expressed as
$\small\begin{split}H\left(Y{{}^{\prime}}\right)-H\left(Y\right)&\leq L_{3}\leq H\left(Y{{}^{\prime}}\right).\end{split}$ (26)
$\small\begin{split}\rm{Thus},~H\left(Z{{}^{\prime}}\right)\!-\!H\left(Y{{}^{\prime}}\right)&\leq L_{2}\!-\!L_{3}\leq H\left(Z{{}^{\prime}}\right)\!-\!H\left(Y{{}^{\prime}}\right).\end{split}$ (27)
Due to Sandwich Theorem and condition A, we have $L_{2}-L_{3}=0$, such that
$\small\begin{split}L&\leq I\left({Z,Z{{}^{\prime}}}\right)-\left[H\left(Y{{}^{\prime}}\right)-H\left({Y{{}^{\prime}}|Y}\right)\right]\\ &=I\left({Z,Z{{}^{\prime}}}\right)-I\left({Y{{}^{\prime}},Y}\right).\end{split}$ (28)